OpenStudy (anonymous):

Solve. n + (–14) = 9

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

well, you gave me a medal fast lol thanks :)

OpenStudy (bass53):

Simplifying n + -9 = -14 Reorder the terms: -9 + n = -14 Solving -9 + n = -14 Solving for variable 'n'. Move all terms containing n to the left, all other terms to the right. Add '9' to each side of the equation. -9 + 9 + n = -14 + 9 Combine like terms: -9 + 9 = 0 0 + n = -14 + 9 n = -14 + 9 Combine like terms: -14 + 9 = -5 n = -5 Simplifying n = -5

OpenStudy (bass53):

is that better @PaulaLovesSchool13

OpenStudy (anonymous):

Yes, thank you bass! Medal for you! :)

OpenStudy (bass53):

yay lol